OpenStudy (anonymous):

A one month supply of multivitamins cost $28 . A two month supply cost $49 . How much money is a guest saving by purchasing a two month supply ?

OpenStudy (nincompoop):

\[\frac{ 49 }{ 28 \times 2 }\times 100\]

OpenStudy (mathstudent55):

@nincompoop Nothing about percentages is mentioned here.

OpenStudy (mathstudent55):

How much would it cost to buy one month's worth and then another month's worth? How much does it cost to buy the two month's worth in one purchase? Subtract the second amount from the first amount.

OpenStudy (anonymous):

i dont understand @mathstudent55 and @nincompoop